Reduce the following fractions to the simplest form: 
 48 / 60 

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

48 / 60 = (12× 4) / (12 × 5)
= 4 / 5

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
Simplest form of the following fraction is 
48/60
= (48÷12) / (60÷12)
= 4/5
 
Hope it helps!
Thank you ✌️